A 1200-kg car going 30 m/s applies its brakes and skids to rest. If the friction force between the sliding tires and the pavement is 6000 N, how far does the car skid before coming to rest?
Can you help me get the formula to use .
Thanks
Answer #1
Set the work done against friction equal to the initial kinetic energy and solve for X

(1/2) M Vo^2 = 6000 * X

X = (1/1200)M Vo^2

Vo is the initial velocity, 30 m/s.
answered by: Thomas S.